OpenStudy (anonymous):

A spring requires a force of 25 N to produce an extension of 2 cm in it. What is the work done in further extending it by 3 cm?

OpenStudy (anonymous):

F= ke use it to find spring constant(k) .can u?

OpenStudy (anonymous):

???

OpenStudy (anonymous):

25= k*2/100 k=1250 Now, work done =1/2 ke^2 =1/2 * 1250 *(5/100)^2

OpenStudy (anonymous):

does it help?

OpenStudy (anonymous):

yep i got it